Normally Sarita takes 3 hours to travel between two stations with a constant speed.One day her speed was reduced by 12 km/hr,and she took 75 minutes more to complete the journey.Calculate the distance between the two stations.
Answer by mananth(16949)

let speed be x km/h
distance = speed * time
distance = 3x
speed reduced by 12 km/h
so speed = x-12 km/h
time = 3h + 1h 15 minutes===>4 1/4 h
17/4 h
17/4 *( x-12)= 3x
17(x-12)=4*3x
17x-204 = 12x
17x-12x=204
5x=204
x= 40.8 km/h
distance = 3*40.8= 122.4km
m.ananth@hotmail.ca